# Towards Generalised Pre-Training of Graph Models

### v3

@@ 4.3 Transfer
Node & Edge Transfer By excluding the output head (again see Figure 1) we are able to transfer ToP pre-trained models onto node and edge level tasks. Here we focus on edge prediction and node classification, and though ToP pre-training does seem to offer some performance increase, the benefits are far smaller than on graph level tasks. These results can be found in the Supplemental Material in Table 10.
@@
Further results can be found in the Supplemental Material in Section B: see Table 5 for transfer with a linear models, Tables (6,7) for GraphCL-like augmentations, Table 8 for a limited ablation over non-molecular pre-training datasets, Table 9 for ToP with varied GNN backbones, and Table 10 for node and edge classification. A comparison of ToP against domain-specific methods is in Table 11.
@@ 5.1 Findings
Pre-training on multiple domains of topologies offers significant performance improvements on downstream tasks (Hypothesis 1). With features excluded we find entirely positive transfer across a range of domains and graph-level tasks using adversarially pre-trained models with ToP. With features included, on popular molecular benchmarks, there are significant performance increases on graph-level tasks compared to fully supervised models.
@@
The most intuitive assumption would be that exposing a molecular representation learner to non-molecular graphs would hinder its performance. Instead the ToP-Social and ToP-All models out-perform the ToP-Chem model almost all molecular datasets under fine-tuning. This holds true for varied pre-training regimes and model backbones (see Tables 67 and 9 in the Supplemental Material). This shows potential for future work using ToP. Pre-training on graphs could be possible through data scale and diversity in-general, instead of within a specific domain. This has clear parallels to the current trends in generalised pre-training in other fields [23, 33].
@@
Similarly one might expect that including node and edge features during fine-tuning (or supervised training of a non-pretrained model) would likely lead to these domain-specific models out-performing all-domain models. Our results instead show that performance is both superior and more consistent from our all-domain model on featured molecular graphs, with performance increasing compared to the same model on non-featured graphs. This allows models, after using ToP, to be transferred to arbitrary tasks on graph structures, without the inherent costs of LLM-based approaches [19].
@@
Given these findings the reasonable use-cases for models pre-trained with ToP are broad. The potential benefit is two-fold. Firstly, pre-training and fine-tuning requires less data for the downstream task to be viable. This is particularly valuable in areas of data scarcity or expense, as well as offering presumed decreases in training time for the final model. Secondly, as shown by our results in this work, our pre-trained models offer actual performance increase over simple supervised training. Our assumption is that the broad range of generic topological features learnt by the ToP model provide a more robust foundation for training than simple random initialisation, so pre-trained models avoid overfitting to non-useful features better than randomly initialised models. Guarantees on this point are left for future research.
5.2 Assumptions, Risks, Future Directions
The most significant assumption we made — that a model can learn representations that are useful on multiple domains of downstream tasks — has been justified by our results. However, in the full scope of downstream tasks for graph data, this first exploratory work is still fairly limited. Firstly we do not experiment with novel graph augmentations (or “views”), instead employing only node and edge dropping. While we’ve shown that adversarial augmentations do improve pre-training, adversarial augmentations could be developed that include some techniques present in Hassani and Khasahmadi [13], for example edge addition alongside edge dropping. Recent works have found that such augmentations, including combined additive and subtractive edge-based augmentations, can lead to more robust representation learners [9]. We anticipate that a variety of augmentations, paired with a more detailed study of how different compositions of domain data influence representations, could bolster the expressivity of these topology-pre-trained models.
Secondly our experiments are limited to a few set compositions of domains. The non-task-specific evaluation of representation models is a challenging and open research problem. The same challenge is compounded on the high-dimensional graph space, which in comparison to other data structures both lacks intuitive human understanding and has a smaller body of existing research. The usual workaround for this issue is increasing the scale and breadth of the pre-training data [28]. Future work using ToP is likely to follow this same path.
Additionally this work primarily focusses on graph-level tasks, in the same manner as the original AD-GCL work [32]. We anticipate that future work will optimise ToP, probably through targeted augmentation schemes, for node and edge related tasks specifically.
Lastly we should note that we haven’t compared ToP to other graph pre-training methods. Such methods are, with a few exceptions [19], domain specific. As such comparisons would be between our generalised models and models pre-trained on a given domain with all features included during pre-training. Such pre-training requires a large dataset of labelled samples, significantly limiting viable domains. The Supplemental Material contains a comparison of ToP pre-trained models with AD-GCL [32], GraphCL [13] and InfoGraph [36].
@@ 5.2 Assumptions, Risks, Future Directions
Liu et al. [21] recently proposed definitions for, and expected properties of, graph foundation models. The models produced with ToP satisfy their broad definition, ‘…benefit from the pre-training of broad graph data, and can be adapted to a wide range of downstream graph tasks’. The two further properties proposed by Liu et al. [21], emergence and homogenization, are less easy to demonstrate. Liu et al. themselves note that how these properties manifest is an open question, and therefore we similarly leave whether ToP models exhibit these properties as an area for future research.
@@ 6 Conclusion
We presented ToP, a method for generalised multi-domain pre-training of graph models. ToP relies on excluding node and edge features during pre-training. Features can then be re-included during fine-tuning on downstream tasks.
Findings Through contrastive learning we used ToP to present graph models capable of positive transfer across domains and tasks. These are an important step towards graph foundation models with use-cases in the manner of models such as BERT [6]. Like BERT, an early foundation language model, ToP models can be fine-tuned with positive transfer on multiple domains and tasks. On 75% of downstream tasks from multiple domains we show significant ($p\leq 0.01$) positive transfer compared to a non-pre-trained model with the same architecture, as expected from our first hypothesis. This is also true when node and edge features are re-introduced during fine-tuning. Compared to a supervised baseline, where results are not significantly better, they are never significantly worse.
Additionally we showed that ToP pre-training on a single target domain leads to at-best on-par, and on most tasks significantly worse ($p\leq 0.01$), performance compared to pre-training on multiple non-target domains. This clearly demonstrates that the variety provided by out-of-domain pre-training can hold more value than the status-quo in-domain pre-training, as in our second hypothesis.
Directions We anticipate that future work will focus on several key areas: 1) Improved or tailored graph augmentation strategies. 2) Investigation of domain weighting in dataset compositions. 3) A larger pre-training dataset spanning a wider range of domains. 4) Improved performance on node-level and edge-level tasks.
Appendix A Supplemental Material
A.1 Dataset Statistics
Table 3: Statistics for each of our datasets. If used for training, we denote these statistics averaged across the training data. Otherwise we report statistics for their validation (downstream fine-tuning) datasets. ‘BC’ indicates binary classification.
@@ A.1 Dataset Statistics
Twitch 50000 29.6 $\pm$ 11.1 86.6 $\pm$ 70.7 2$\pm$ 0 0.549 $\pm$ 0.149 BC. Games Played
@@
molclintox 1329 26.3 $\pm$ 15.8 28.0 $\pm$ 17.1 12.4 $\pm$ 6.09 0.00259 $\pm$ 0.0191 BC. Toxicity
molbbbp 1835 23.5 $\pm$ 9.89 25.4 $\pm$ 11.0 11.2 $\pm$ 4.03 0.00284 $\pm$ 0.0278 BC. Penetration
molbace 1361 34.0 $\pm$ 7.89 36.8 $\pm$ 8.13 15.2 $\pm$ 3.14 0.00672 $\pm$ 0.0205 BC. Bioactivity
molhiv 37014 25.3 $\pm$ 12.0 27.3 $\pm$ 13.1 12.0 $\pm$ 5.15 0.00158 $\pm$ 0.0156 BC. Inhibition
@@
A.2 Exploration of Embedding Space
Figure 3: A UMAP embedding of encodings from each model, as well as an untrained model. The untrained and Chem models are noticeably more fragmented than the Social and All models. In turn the Social model is more fragmented than the All model. In the All model embedding, molecules (primarily dark blue) are all in the same region, furthest from the Twitch Ego networks (lilac), with the Cora (brown) and Facebook Page-Page data (pink) “filling the gap”.
On the combined encodings of the validation sets we compute $R^{2}$ correlation coefficients between the first five principal components (PCA) and common graph-level metrics. High correlations, or the lack thereof, should give some indication of how embedding components represent different graph characteristics. The results are presented, along with each component’s explained variance, in Table 4 and visualised in Figure 4.
Table 4: PCA components, fit on the embeddings of the test set, and their most correlated graph-level metric, for each of the trained models. Underlined text indicates a strong correlation $|R^{2}|\geq 0.66$, and bold text indicates moderate correlation $0.33\leq|R^{2}|\leq 0.66$. We show up to the first five PCA components, but include only three where explained variance values are minimal.

### v4

@@ 4.3.3 Further Transfer Results
By excluding the output head (again see Figure 2) we are able to transfer ToP pre-trained models onto node and edge level tasks. Here we focus on edge prediction and node classification, and though ToP pre-training does seem to offer some performance increase, the benefits are far smaller than on graph level tasks. These results can be found in the Appendix in Table 12.
@@
Further results can be found in the Appendix in Section C: see Table 8 for transfer with a linear models, Tables (9,10) for random GraphCL-like augmentations, Table 11 for ToP with varied GNN backbones, and Table 12 for node and edge classification.
4.4 Further Experiments
Having demonstrated the efficacy of ToP, we present two more experiments, designed to give some insight into why and how ToP pre-training is effective. The first ensures that our results supporting Hypothesis 2, that out-of-domain data can be more useful than in-domain, are not the result of artefacting. The second investigates the extent to which ToP, during downstream transfer, relies on feature information in comparison to structural information.
4.4.1 Varied Pre-Training Data
While the results we present are statistically significant, there remains the possibility of artefacting. In particular it is feasible that a particular domain in the ToP-Social pre-training data lead to positive transfer, not the overall diversity of data. In order to study and eliminate that possibility we perform a limited ablation over the component dataset within the Social pre-training data. We exclude one of the five pre-training datasets at a time, pre-train an encoder with ToP, then fine-tune ten times across our validation datasets, under the same transfer settings as our main results. These results are presented in Table 5.
Table 5: Scores for fine-tuning models pre-trained on non-molecules, each with one non-molecular pre-training dataset excluded. For the classification datasets we report AUROC and for regression datasets RMSE. Bold text indicates a result performs equally (or better) than the model pre-trained on molecules (Chem). Molecular tasks include domain features in fine-tuning.
Social (Name indicates excluded dataset)
Twitch Neurons Cora Facebook Roads
Facebook 0.29 $\pm$ 0.02 0.14 $\pm$ 0.00 0.84 $\pm$ 0.01 0.14 $\pm$ 0.01 0.22 $\pm$ 0.06
Twitch Egos 0.73 $\pm$ 0.03 0.74 $\pm$ 0.01 0.73 $\pm$ 0.01 0.74 $\pm$ 0.01 0.77 $\pm$ 0.01
Roads 0.19 $\pm$ 0.02 0.11 $\pm$ 0.00 0.12 $\pm$ 0.01 0.11 $\pm$ 0.00 0.18 $\pm$ 0.01
Trees 0.15 $\pm$ 0.01 0.12 $\pm$ 0.00 0.13 $\pm$ 0.01 0.12 $\pm$ 0.00 0.15 $\pm$ 0.01
Community 0.28 $\pm$ 0.03 0.02 $\pm$ 0.00 0.07 $\pm$ 0.01 0.02 $\pm$ 0.01 0.22 $\pm$ 0.03
Random 0.15 $\pm$ 0.01 0.05 $\pm$ 0.00 0.09 $\pm$ 0.02 0.04 $\pm$ 0.00 0.13 $\pm$ 0.02
molfreesolv 3.30 $\pm$ 0.32 3.90 $\pm$ 0.24 3.82 $\pm$ 0.31 3.67 $\pm$ 0.37 3.83 $\pm$ 0.38
molesol 1.03 $\pm$ 0.08 1.03 $\pm$ 0.07 1.26 $\pm$ 0.16 1.13 $\pm$ 0.16 1.17 $\pm$ 0.10
mollipo 0.93 $\pm$ 0.03 0.83 $\pm$ 0.03 0.92 $\pm$ 0.05 0.81 $\pm$ 0.03 0.94 $\pm$ 0.02
molclintox 0.49 $\pm$ 0.03 0.76 $\pm$ 0.13 0.72 $\pm$ 0.10 0.76 $\pm$ 0.18 0.74 $\pm$ 0.06
molbbbp 0.64 $\pm$ 0.02 0.64 $\pm$ 0.02 0.61 $\pm$ 0.03 0.62 $\pm$ 0.02 0.64 $\pm$ 0.03
molbace 0.70 $\pm$ 0.04 0.69 $\pm$ 0.07 0.68 $\pm$ 0.05 0.67 $\pm$ 0.06 0.72 $\pm$ 0.05
molhiv 0.41 $\pm$ 0.06 0.55 $\pm$ 0.11 0.48 $\pm$ 0.13 0.63 $\pm$ 0.07 0.52 $\pm$ 0.11
We find that while performance does in-general drop compared to ToP-Social, it is still consistently higher than ToP-Chem. Some datasets do appear to have a larger benefit for transfer than others - for example, when the Twitch Ego networks are excluded, performance on the molclintox dataset drops to near-random. It is worth noting that we do not scale-up the Social pre-training data when a dataset is included, so these ablation models are all pre-trained on 20% less data than ToP-Chem and ToP-Social.
4.4.2 Features vs. Structure
Assume some imbalance between features and structure, $I(y;X)\neq I(y;E)$. Here we propose a way of investigating this imbalance between feature information and structure information. By degrading the useful information in either (or both of) $X$ and $E$, the degree to which performance relies on one or the other should be apparent. A caveat here is that degrading either will presumably also degrade the information from their interdependence.
Consider some destructive noising process $N_{t}(\cdot)$, that degrades the useful information in either $X$ or $E$. $N_{0}(x)=x$, with no useful information after $N_{T}(\cdot)$.
From here edge noising is denoted $N_{E}=N_{t}(E)$, and node feature noising $N_{X}=N_{t}(X)$ Assume some imperfect model produces predictions $f(X,E)\rightarrow\hat{y}$. Performance is some $h(y,\hat{y})$ between real and predicted values. By sampling $h(y,f(N_{X},N_{E})\rightarrow\hat{y})$, the useful information in both $X$ and $E$, and whether performance requires both to be present, should be observable. We apply this experiment over our molecular benchmarks. Here $X$ includes both node and edge features.
Structure noise
We use random edge removal/addition for structure noise. At $t$, we remove $|E_{r}|=p_{t}\cdot|V|$ edges at random from the graph, $E^{\prime}=E\setminus E_{r}$. We then add an equal number of randomly chosen edges $|E_{a}|=p_{t}\cdot|V|$, for a final noised edgelist $E^{\prime\prime}=E^{\prime}\cup E_{a}$.
This is very close to the procedure taken by graph diffusion models as forward noising processes [46]. At $N_{T}(E)$, the structure is a random graph with the same density as the original.
Feature Noise - Permutation
We employ additive Bernoulli noise across node features. At $t$, we have a probability of permuting the entries across a feature $p_{t}$. Let node (or edge features) across a whole dataset be $X\in\{0,1\}^{N\times D}$ with $N$ points and $D$ features. At $t$ we randomly select $N_{f}=p_{t}\cdot D$ features to permute. Each selected feature is then shuffled, across the whole dataset, removing the useful information from that feature that pertains to each sample, while maintaining the marginal distribution of the given feature.
Results
We fix the test set of each benchmark as un-noised, then perform 8 supervised training runs across 8 increasing noise levels on features and structure. In contrast to our transfer results we perform these supervised training runs only for 25 epochs. When noise is applied to structure, no noise is applied to features, and vice-versa. We present the results in Figure 6 for regression datasets and Figure 10 for classification datasets.
Performance suffers by varying amounts across different datasets. On most datasets the untrained GIN’s response to structural noise is minimal, with little performance change even at complete noise in the graph structure. The untrained GIN’s reponse to feature noise is far larger, in particular on regression datasets (molesol, molfreesolv, mollipo). The Chem model, pre-trained only on molecules, follows the same patterns as the untrained GIN model.
On most datasets the two non-molecular models, Social and All, show large performance drops as both structural and feature noise increases. This is most clear on the datasets where ToP pre-training was most effective, for example molclintox. Interestingly this suggests that ToP pre-training allows models not only to use structural patterns to greater effect, but also integrate information from features more usefully. An additional possibility is that ToP pre-training protects against overfitting to patterns in graph features, hence those features being used to greater effect.
Figure 6: Performance variation for each molecular regression benchmark dataset with increasing noise on structure and features. For each dataset, we benchmark performance for an untrained GIN and our three models pre-trained with ToP (Chem-100, Social-100, All-100).
Figure 10: Performance variation for each molecular classification benchmark dataset with increasing noise on structure and features. For each dataset, we benchmark performance for an untrained GIN and our three models pre-trained with ToP (Chem-100, Social-100, All-100).
@@ 5.1 Findings
Pre-training on multiple domains of topologies offers significant performance improvements on downstream tasks (Hypothesis 1). With features excluded we find entirely positive transfer across a range of domains and graph-level tasks using adversarially pre-trained models with ToP. With features included, on popular molecular benchmarks, there are significant performance increases on graph-level tasks compared to fully supervised models. This means, broadly, that ToP can be used without risk of negative transfer.
@@ 5.2 Intuitions and Comparisons
The most intuitive assumption would be that exposing a molecular representation learner to non-molecular graphs would hinder its performance. Instead the ToP-Social and ToP-All models out-perform the ToP-Chem model on almost all molecular datasets under fine-tuning. This holds true for varied pre-training regimes and model backbones (see Tables 910 and 11 in the Appendix). This shows potential for future work using ToP. Pre-training on graphs could be possible through data scale and diversity in-general, instead of within a specific domain. This has clear parallels to the current trends in generalised pre-training in other fields [30, 47].
@@
Similarly one might expect that including node and edge features during supervised training of a non-pretrained model would likely lead to these domain-specific models out-performing our ToP multi-domain models. Our results instead show that performance is both superior and more consistent from our all-domain model on featured molecular graphs, with performance increasing compared to the same model on non-featured graphs.
In comparison to other generalist graph approaches, ToP models perform strongly, including against science-aimed LLM models many orders of magnitude larger in both training data and cost of use. This allows models, after using ToP, to be transferred to arbitrary tasks on graph structures, without the inherent costs of LLM-based approaches.
We extend, using techniques from this work, GCC models to include domain features. Here pre-trained GCC encoders show consistently negative transfer. This indicates that GCC’s use of positional encodings as node and edge features during pre-training actively inhibits their ability to use the information from domain features, which in most applications of graph models are essential for strong performance. ToP models, in contrast, show large performance increases when domain features are included during downstream fine-tuning compared to when domain features are not included. Positive transfer is at least as common with these features included as when they are not.
@@
Given these findings the reasonable use-cases for models pre-trained with ToP are broad. The potential benefit is two-fold. Firstly, pre-training and fine-tuning requires less data for the downstream task to be viable. This is particularly valuable in areas of data scarcity or expense, as well as offering presumed decreases in training time for the final model. Secondly, as shown by our results in this work, our pre-trained models offer actual performance increase over simple supervised training. Our assumption is that ToP pre-training introduces a strong bias towards structural information, which in many applications is a positive characteristic, in-particular where it inhibits overfitting to patterns in the space of node or edge features. Guarantees on this point are left for future research.
Graph Foundation Models
@@
Liu et al [26] recently proposed definitions for, and expected properties of, graph foundation models. The models produced with ToP satisfy their broad definition, ‘…benefit from the pre-training of broad graph data, and can be adapted to a wide range of downstream graph tasks’. The two further properties proposed by Liu et al [26], emergence and homogenization, are less easy to demonstrate. Liu et al.themselves note that how these properties manifest is an open question, and therefore we similarly leave whether ToP models exhibit these properties as an area for future research.
A useful lens is to compare ToP to the much more mature field of foundation models for natural language, ie LLMs. While the current batch of LLMs are deeply expressive, and capable of performing a massive range of natural language tasks at a high level, earlier iterations of the same fundamental technology were more rudimentary. The best parallel for use-cases of ToP might be BERT [9]. BERT is, by current standards, quite a small LLM, and normally requires fine-tuning to reach good performance on downstream tasks. We’re confident that future graph foundation models will be both more powerful and more flexible in terms of tasks, in particular zero-shot and few-shot learning, and leave whether ToP models are an early version of those GFMs as an open question.
5.3 Limitations
In the full scope of pre-training and downstream tasks for graph data, this work has some limitations. Firstly we use only static, non-hyper graphs, as in many other graph learning works. We do not experiment with novel graph augmentations (or “views”), instead employing only node and edge dropping, both random and adversarial. We show that ToP offers performance benefits with random views, and that adversarial views bolster these performance benefits. While adversarial augmentations do improve pre-training, novel augmentations could be developed that include some techniques present in Hassani and Khasahmadi [17], for example, edge addition alongside edge dropping. Recent works find that such augmentations, including combined additive and subtractive edge-based augmentations, can lead to more robust representation learners [13].
Our method primarily focuses on graph-level tasks, in the same manner as the original AD-GCL work [43]. As such, and again as in that original work, we find that performance benefits on node or edge level tasks are much smaller than on graph-level tasks. For node and edge related tasks, future works can optimise ToP through targeted augmentation schemes. Similarly, we employ a limited set of domain compositions in evaluating ToP. A variety of augmentations, paired with a more detailed study of how different compositions of domain data influence representations, could bolster the expressivity of these topology-pre-trained models.
@@ 6 Conclusion
We present ToP, a method for generalised multi-domain pre-training of graph models. ToP relies on excluding node and edge features during pre-training. Features can then be re-included during fine-tuning on downstream tasks.
Through contrastive learning, we use ToP to present graph models capable of positive transfer across domains and tasks. These are an important step towards graph foundation models with use-cases in the manner of models such as BERT [9]. Like BERT, an early foundation language model, ToP models can be fine-tuned with positive transfer on multiple domains and tasks. On 75% of downstream tasks from multiple domains we show significant ($p\leq 0.01$) positive transfer compared to a non-pre-trained model with the same architecture, as expected from our first hypothesis. This is also true when node and edge features are re-introduced during fine-tuning. Compared to a supervised baseline, where results are not significantly better, they are never significantly worse. Additionally we showed that ToP pre-training on a single target domain leads to at-best on-par, and on most tasks significantly worse ($p\leq 0.01$), performance compared to pre-training on multiple non-target domains. This clearly demonstrates that the variety provided by out-of-domain pre-training can in this setting hold more value than in-domain pre-training, as in our second hypothesis.
We anticipate that future work will focus on several key areas: 1) Improved or tailored graph augmentation strategies. 2) Investigation of domain weighting in dataset compositions. 3) A larger pre-training dataset spanning a wider range of domains. 4) Improved performance on node-level and edge-level tasks. 5) Two-stage pre-training with an in-domain, features-included method. 6) Integration and adaption for graph transformers
[…]
Supplementary information
All code for ToP, including data processing, result-reproduction and pre-trained models is available in our GitHub repository https://github.com/alexodavies/general-gcl. Our updated code for GCC, including the inclusion of features in downstream tasks, is available in its own GitHub repository https://github.com/alexodavies/GCC.
[…]
Acknowledgements
AOD and RG acknowledge support from the UK Research and Innovation (UKRI) Centre for Doctoral Training in Interactive Artificial Intelligence Award (EP/S022937/1).
Appendix A Dataset Statistics
Table 6: Statistics for each of our datasets. If used for training, we denote these statistics averaged across the training data. Otherwise we report statistics for their validation (downstream fine-tuning) datasets. All tasks are single-target.
@@ Appendix A Dataset Statistics
Twitch 50000 29.6 $\pm$ 11.1 86.6 $\pm$ 70.7 2$\pm$ 0 0.549 $\pm$ 0.149 Class. Games Played
@@
molclintox 1329 26.3 $\pm$ 15.8 28.0 $\pm$ 17.1 12.4 $\pm$ 6.09 0.00259 $\pm$ 0.0191 Class. Toxicity
molbbbp 1835 23.5 $\pm$ 9.89 25.4 $\pm$ 11.0 11.2 $\pm$ 4.03 0.00284 $\pm$ 0.0278 Class. Penetration
molbace 1361 34.0 $\pm$ 7.89 36.8 $\pm$ 8.13 15.2 $\pm$ 3.14 0.00672 $\pm$ 0.0205 Class. Bioactivity
molhiv 37014 25.3 $\pm$ 12.0 27.3 $\pm$ 13.1 12.0 $\pm$ 5.15 0.00158 $\pm$ 0.0156 Class. Inhibition
@@
Appendix B Embedding Space
Figure 11: A UMAP embedding of encodings from an untrained encoder.
Figure 12: A UMAP embedding of encodings from the ToP-Chem model, pre-trained on only molecules.
Figure 13: A UMAP embedding of encodings from the ToP-Social model, pre-trained on only non-molecules.
Figure 14: A UMAP embedding of encodings from the ToP-All model, pre-trained on the whole amalgamated dataset (i.e. molecules and non-molecules).
Table 7: On the combined encodings of the validation sets we compute $R^{2}$ correlation coefficients between the first five principal components (PCA) and common graph-level metrics. High correlations, or the lack thereof, should give some indication of how embedding components represent different graph characteristics. Here we present the most correlated metrics and components for each pre-trained model. Underlined text indicates a strong correlation $|R^{2}|\geq 0.66$, and bold text indicates moderate correlation $0.33\leq|R^{2}|\leq 0.66$. We show up to the first five PCA components, but include only three where explained variance values are minimal.